OpenStudy (anonymous):

Find the slope of the line passing through the points (-2, 9) and (5, 3). A. -6/7 B. -7/6 C. 4 D. 1/4

OpenStudy (anonymous):

@texaschic101 I say -7/6 but wouldnt u change it to -6/7?

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (-2,9)...x1 = -2 and y1 = 9 (5,3)...x2 = 5 and y2 = 3 now we sub slope = (3 - 9) / (5 - (-2) slope = -6/(5 + 2) slope = -6/7

OpenStudy (anonymous):

Find the slope of line passing through the points (4,8) and (6, -2). I think is -1/5 @texaschic101